Jessicakuck:

total amount = P (1 + i)t What is the total amount that First Consumer Bank will receive after lending Jane $7,000 for three years at an interest rate of 5 percent, compounded annually? $8,103.38 $8,114.92 $8,358.37

Jessicakuck:

@TylerJoseph

Jessicakuck:

@SkyVoltage43

21Pilots:

A

TylerJoseph:

Tyler is correct. It would be A. :)

SkyVoltage43:

yep a

Jessicakuck:

simple interest = P × r × t Dave has a three–year college loan for $10,000. The lender charges a simple interest rate of 6 percent. How much interest will he have to pay?

SkyVoltage43:

1.If Dave has a three-year college loan for $10,000 and the lender charges a simple interest rate of 6 percent, Dave will have to pay $1,800 in interest. $10,000 principle at 6 percent annual interest rate is equal to a monthly interest payment of $50. 36 X $50 = $1,800.

Jessicakuck:

total amount = P (1 + i)t Ryan has an eight–year loan for $6,000. He is being charged an interest rate of 5 percent, compounded annually. Calculate the total amount that he will pay. Brainly isn't working. sorry guys hahaha

SkyVoltage43:

Given: Principal : $6,000 Interest Rate: 5% Term : 8 years, compounded annually. The term compounded annually is a hint that informs us to use the compounded interest formula instead of the simple interest formula. Compounded interest formula is: A = P(1 + r/n)^nt where: A = future value of loan or investment including the interest P = principal r = rate n = the number of times the interest is compounded per year t = the number of years the money is borrowed or invested A = P (1 + r/n)^nt A = 6,000 (1 + 0.05/1)¹ * ⁸ A = 6,000 (1.05)⁸ A = 6,000 (1.48) A = 8,880 The total amount Ryan will pay after 8 years is $8,880.00

Jessicakuck:

That isn't one of the options thoughh

Jessicakuck:

$7,657.69 $8,815.97 $8,864.73

SkyVoltage43:

hmmm did i math wrong

SkyVoltage43:

its C my bad

Jessicakuck:

I guessed.. and failed that test.

SkyVoltage43:

O-O